OpenStudy (anonymous):

Find the difference. Write your answer in simplest form. 1 1/6 - 5/9

OpenStudy (anonymous):

1 1/6 - 5/9 = 11/18 in simplest form

OpenStudy (anonymous):

@anniedili100

OpenStudy (irishboy123):

when you are using rational (fractional) numbers, you need to start lining up the numbers on the bottom of the fraction, viz the denominators, but first in this case sort out the top. 1 1/6 is 7/6 {ie 6/6 + 1/6 = 7/6} so, you are finding the difference between 7/6 and 5/9. you need to get the numbers on the bottom to agree or you are comparing apples to pears and the answer is meaningless. KEY POINT: 6 * 9 = 54. 7/6 = 63/54 5/9 = 30/54 ergo: (63-30)/54 = 33/54 = 11/18